OpenStudy (anonymous):

Condense the Logarithmic Expression: ln12-2lnx

OpenStudy (mathmale):

Three important rules of logs include the following: 1) ln a + ln b = ln a*b 2) ln a- ln b = ln (a/b) 3) ln a^b = b*ln a Use #3 to change 2 ln x to its alternative form.